A 1 kg object is located at a distance of 6.4x106 m from the center of a larger object whose mass is 6.0x1024 kg.

a. What is the size of the force acting on the smaller object?

 b. What is the size of the force acting on the larger object?


1
Expert's answer
2021-12-20T10:32:00-0500

The gravity force (regardless of the object) is given as follows:


"F = G\dfrac{m_1m_2}{r^2}"

where "G = 6.8 \u00d7 10^{-11} m^3 kg^{-1} c^{-2}", "m_1 = 1kg, m_2 = 6.0\times 10^{24}kg, r = 6.4\times 10^6m". Thus, obtain:


"F = 6.8\times 10^{-11}\cdot \dfrac{1\cdot 6\times 10^{24}}{(6.4\times 10^6)^2} \approx 10N"

Answer. a) 10N, b) The same, 10N.